# Diffraction by a Right-Angle Penetrable Wedge: Closed-Form Solution for $\nu=\sqrt{2}$

### v2

@@ 3 Weierstrass uniformization for $\tau=i$
Let $\wp(u)$, $\wp^{\prime}(u)$ and $\zeta_{W}(u)$ denote the corresponding Weierstrass elliptic and zeta functions; see, e.g., [14, 15, 16]. We adopt the uniformization
$\wp(u)=x_{W}(t):=\frac{it^{2}+1}{it^{2}-1},\qquad\frac{1}{2}\,\wp^{\prime}(u)=y_{W}(t,Y):=\frac{(i-1)tY}{(it^{2}-1)^{2}}.$ (18)
The physical lift $u=u(\zeta)$ is selected by composing the physical branch $\zeta\mapsto(t(\zeta),Y(\zeta))$ with the uniformization on $\Omega_{+}^{\mathrm{phys}}$.
Let $u_{0}$ denote the half-period corresponding to the physical point $(t,Y)=(0,-\sqrt{2})$, so that
@@ 4 Pole set, labels, and incident exclusion
$(+,+)\mapsto 3,\qquad(+,-)\mapsto 1,\qquad(-,+)\mapsto 4,\qquad(-,-)\mapsto 2,$ (21)
@@
For later reference we make the pole label $\ell\mapsto(t_{\ell},Y_{\ell})\mapsto u_{\ell}$ explicit. Set the limiting-absorption incident angle $\zeta_{i}=\theta_{i}+i\varepsilon$ and fix $\theta_{w}=\pi/4$. For $(\sigma,\varepsilon_{w})\in\{\pm 1\}^{2}$ define the four forcing phases
$a_{\sigma,\varepsilon_{w}}:=\exp\bigl(i\sigma(\zeta_{i}+\varepsilon_{w}\theta_{w})\bigr),\qquad b_{m,\sigma,\varepsilon_{w}}:=a_{\sigma,\varepsilon_{w}}^{(-1)^{m}}\quad(m=0,1,2,3).$ (24)
The physical orbit table fixes the pole condition in the form
$e^{iw_{m}}=(s(\tau^{m}p))^{(-1)^{m}}=b_{m,\sigma,\varepsilon_{w}},$
so that the forcing pole points are the solutions of $s(q)=b$ on $\Sigma_{\mathrm{lem}}$ with $q=\tau^{m}p$.
@@
On the physical component $\Omega_{+}^{\mathrm{phys}}$ one selects the root $t_{\mathrm{in}}(b)\in\mathbb{D}:=\{|t|<1\}$, and the other root is $t_{\mathrm{out}}(b)=1/t_{\mathrm{in}}(b)$.
@@
$\wp(u_{\ell})=x_{W}(t_{\ell}),\qquad\frac{1}{2}\,\wp^{\prime}(u_{\ell})=y_{W}(t_{\ell},Y_{\ell}),$ (29)
with the physical lift selected on $\Omega_{+}^{\mathrm{phys}}$ (equivalently, by fixing the sign of $\wp^{\prime}(u_{\ell})$ through $y_{W}(t_{\ell},Y_{\ell})$).
@@ 5.1 Derivative and residue conventions
$w_{m}^{\prime}(u)=\frac{\,\mathrm{d}w_{m}}{\,\mathrm{d}u}=\begin{cases}-e^{-i\pi/4}\Bigl(t-\dfrac{1}{t}\Bigr),&m\ \text{even},\\[4.30554pt] +e^{-i\pi/4}\Bigl(t+\dfrac{1}{t}\Bigr),&m\ \text{odd},\end{cases}\qquad t=t(u).$ (30)
@@
The orbit derivatives in (30) follow from the lemniscatic Snell relation (13) and the uniformization (18).
@@
Write $w=w_{0}$ for the physical branch defined by $e^{iw}=s(t,Y)$, and set $t=t(u)$, $Y=Y(u)$ along the physical lift. Differentiating (18) gives
$\wp^{\prime}(u)=\frac{d}{du}x_{W}\bigl(t(u)\bigr)=x_{W}^{\prime}(t)\,t^{\prime}(u).$
A direct computation from $x_{W}(t)=(it^{2}+1)/(it^{2}-1)$ yields
$x_{W}^{\prime}(t)=-\frac{4it}{(it^{2}-1)^{2}},$
while $\wp^{\prime}(u)=2y_{W}(t,Y)=2(i-1)tY/(it^{2}-1)^{2}$. Equating these expressions shows
$\frac{dt}{du}=t^{\prime}(u)=-\frac{1+i}{2}\,Y(u).$ (31)
Next, differentiating the Snell identity $\cos w=\nu\cos z$ with $\nu=\sqrt{2}$ and using $t=e^{iz}$ gives
$\sin w\,\frac{dw}{dz}=-\sqrt{2}\,\sin z\qquad\Rightarrow\qquad\frac{dw}{dz}=\sqrt{2}\,\frac{\sin z}{\sin w}.$
Expressing $\sin z=(t-1/t)/(2i)$ and $\sin w=(s-1/s)/(2i)$, and using the relation $s=s(t,Y)$, one obtains the algebraic form already recorded in (14):
$\frac{dw}{dz}=g^{\prime}(t,Y)=\sqrt{2}\,\frac{t^{2}-1}{Y}.$
Since $dt/dz=it$, we have
$\frac{dw}{dt}=\frac{dw/dz}{dt/dz}=-i\sqrt{2}\,\frac{t^{2}-1}{tY}.$
Combining this with (31) gives
$\frac{dw}{du}=\frac{dw}{dt}\frac{dt}{du}=-i\sqrt{2}\,\frac{t^{2}-1}{tY}\Bigl(-\frac{1+i}{2}Y\Bigr)=-e^{-i\pi/4}\Bigl(t-\frac{1}{t}\Bigr),$
which is the $m$ even case in (30). For the orbit branches $w_{m}$ defined by (12), the additional factors $\tau^{m}$ and $(-1)^{m}$ change $t$ to $i^{m}t$ and introduce the sign change for odd $m$, yielding the $m$ odd case in (30). ∎
@@
Equivalently, the explicit forms are
$r_{I}(\ell)=\begin{cases}-\varepsilon_{j}\,e^{i\pi/4}\dfrac{t_{\ell}}{t_{\ell}^{2}-1},&m\ \text{even},\\[5.16663pt] +\varepsilon_{j}\,e^{i\pi/4}\dfrac{t_{\ell}}{t_{\ell}^{2}+1},&m\ \text{odd}.\end{cases}$ (33)
@@ 5.2 Phase symbols
$\chi_{m}:=i^{m}\in\{\pm 1\}\quad(m\ \text{even}),\qquad\psi_{m}:=i^{m}\in\{\pm i\},\qquad\kappa_{m}:=i^{m+1}\in\{\pm 1\}\quad(m\ \text{odd}).$ (34)
@@ 6 Residue data and per-pole jet summands
6.1 Residue data: $(\alpha_{\ell})$, $(\beta_{\ell})$, $(C_{\ell})$
@@
$\text{if $m$ is even:}\quad\alpha_{\ell}$ $=\begin{cases}-\chi_{m}\,r_{I}\,t^{2},&j=1,\\ -\chi_{m}\,r_{I}\,(t^{4}-t^{2}+1),&j=2,\\ 0,&j=3,\\ -\chi_{m}\,r_{I}\,\dfrac{iY}{\sqrt{2}}(t^{2}-1),&j=4;\end{cases}$ (35)
$\text{if $m$ is odd:}\quad\alpha_{\ell}$ $=\begin{cases}\psi_{m}\,r_{I}\,t^{4},&j=1,\\ \psi_{m}\,r_{I},&j=2,\\ \kappa_{m}\,r_{I}\,\dfrac{Y}{\sqrt{2}}\,t^{2},&j=3,\\ -\kappa_{m}\,r_{I}\,\dfrac{Y}{\sqrt{2}},&j=4.\end{cases}$ (36)
Proof.
See Appendix A. ∎
@@
$\text{if $m$ is even:}\quad\beta_{\ell}$ $=\begin{cases}\chi_{m}\,r_{I}\,\dfrac{i}{\sqrt{2}}Y\,t^{2},&j=1,\\ \chi_{m}\,r_{I}\,\dfrac{i}{\sqrt{2}}Y,&j=2,\\ -\chi_{m}\,r_{I}\,t^{4},&j=3,\\ -\chi_{m}\,r_{I},&j=4;\end{cases}$ (37)
$\text{if $m$ is odd:}\quad\beta_{\ell}$ $=\begin{cases}0,&j=1,\\ \kappa_{m}\,r_{I}\,\dfrac{1}{\sqrt{2}}Y\,(t^{2}+1),&j=2,\\ \psi_{m}\,r_{I}\,t^{2},&j=3,\\ -\psi_{m}\,r_{I}\,(t^{4}+t^{2}+1),&j=4.\end{cases}$ (38)
Proof.
See Appendix A. ∎
@@
$\text{if $m$ is even:}\quad C_{\ell}$ $=\begin{cases}-\chi_{m}\,\dfrac{r_{I}}{2t^{2}},&j=1,\\[3.87495pt] -\dfrac{r_{I}}{2}\,\dfrac{2t^{4}-2t^{2}+1}{t^{4}},&j=2,\ m=0,\\[5.16663pt] +\dfrac{r_{I}}{2}\,\dfrac{1}{t^{4}},&j=2,\ m=2,\\[3.87495pt] 0,&j=3,\\ 0,&j=4;\end{cases}$ (39)
$\text{if $m$ is odd:}\quad C_{\ell}$ $=\begin{cases}0,&j=1,\\ 0,&j=2,\\[3.44444pt] \dfrac{r_{I}}{2}\Bigl(1+\kappa_{m}\,\dfrac{Y}{\sqrt{2}\,t^{2}}\Bigr),&j=3,\\[5.16663pt] -\dfrac{r_{I}}{2t^{2}}\Bigl(1+\kappa_{m}\,\dfrac{Y}{\sqrt{2}\,t^{2}}\Bigr),&j=4.\end{cases}$ (40)
Proof.
See Appendix A. ∎
@@ 6.2 Per-pole jet summands
with coefficients $p_{n}=\sum_{\ell\in I_{\mathrm{scat}}}p_{n}^{(\ell)}$ and $q_{n}=\sum_{\ell\in I_{\mathrm{scat}}}q_{n}^{(\ell)}$. For later reference we record the per-pole contributions to the jet-killing coefficients in closed form.
In this section we record the explicit coefficient formulas; their derivation is given in Appendix A.
@@
Evaluate at $(t,Y)=(t_{\ell},Y_{\ell})$.
  • $m$ even ($\chi_{m}$).
    $j=1:\$$p_{1}^{(\ell)}=\chi_{m}\,i\,\frac{t}{t^{2}-1},\qquad p_{2}^{(\ell)}=-\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Y}{t^{2}-1},\qquad p_{3}^{(\ell)}=\chi_{m}\,i\,\frac{t^{4}+3}{3t(t^{2}-1)}.$
    $j=2:\$$p_{1}^{(\ell)}=\chi_{m}\,i\,\frac{-t^{4}+t^{2}-1}{t(t^{2}-1)},\qquad p_{2}^{(\ell)}=+\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Y(t^{4}-t^{2}+1)}{t^{2}(t^{2}-1)},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=-\chi_{m}\,i\,\frac{(t^{4}+3)(t^{4}-t^{2}+1)}{3t^{3}(t^{2}-1)}.$
    $j=3:\$$p_{1}^{(\ell)}=0,\qquad p_{2}^{(\ell)}=0,\qquad p_{3}^{(\ell)}=0.$
    $j=4:\$$p_{1}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}\,t},\qquad p_{2}^{(\ell)}=-\chi_{m}\,\frac{t^{4}+1}{t^{2}},\qquad p_{3}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t^{4}+3}{3t^{3}}.$
  • $m$ odd ($\psi_{m},\kappa_{m}$).
    $j=1:\$$p_{1}^{(\ell)}=i\,\psi_{m}\,\frac{t^{3}}{t^{2}+1},\qquad p_{2}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Yt^{2}}{t^{2}+1},\qquad p_{3}^{(\ell)}=i\,\psi_{m}\,\frac{t(t^{4}+3)}{3(t^{2}+1)}.$
    $j=2:\$$p_{1}^{(\ell)}=-i\,\psi_{m}\,\frac{1}{t(t^{2}+1)},\qquad p_{2}^{(\ell)}=+\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Y}{t^{2}(t^{2}+1)},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=-i\,\psi_{m}\,\frac{t^{4}+3}{3t^{3}(t^{2}+1)}.$
    $j=3:\$$p_{1}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Yt}{t^{2}+1},\qquad p_{2}^{(\ell)}=-i\,\kappa_{m}\,\frac{t^{4}+1}{t^{2}+1},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y(t^{4}+3)}{3t(t^{2}+1)}.$
    $j=4:\$$p_{1}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y}{t(t^{2}+1)},\qquad p_{2}^{(\ell)}=-i\,\kappa_{m}\,\frac{t^{4}+1}{t^{2}(t^{2}+1)},$
    $\hskip 45.00006ptp_{3}^{(\ell)}=\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y(t^{4}+3)}{3t^{3}(t^{2}+1)}.$
@@
Evaluate at $(t,Y)=(t_{\ell},Y_{\ell})$.
  • $m$ even ($\chi_{m}$).
    $j=1:\$$q_{1}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t}{t^{2}-1},\qquad q_{2}^{(\ell)}=-\chi_{m}\,\frac{t^{4}+1}{t^{2}-1},\qquad q_{3}^{(\ell)}=\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t^{4}+3}{3t(t^{2}-1)}.$
    $j=2:\$$q_{1}^{(\ell)}=-\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{1}{t(t^{2}-1)},\qquad q_{2}^{(\ell)}=+\chi_{m}\,\frac{t^{4}+1}{t^{2}(t^{2}-1)},$
    $\hskip 45.00006ptq_{3}^{(\ell)}=-\chi_{m}\,\frac{Y}{\sqrt{2}}\,\frac{t^{4}+3}{3t^{3}(t^{2}-1)}.$
    $j=3:\$$q_{1}^{(\ell)}=\chi_{m}\,i\,\frac{t^{3}}{t^{2}-1},\qquad q_{2}^{(\ell)}=-\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Yt^{2}}{t^{2}-1},\qquad q_{3}^{(\ell)}=\chi_{m}\,i\,\frac{t(t^{4}+3)}{3(t^{2}-1)}.$
    $j=4:\$$q_{1}^{(\ell)}=-\chi_{m}\,i\,\frac{1}{t(t^{2}-1)},\qquad q_{2}^{(\ell)}=+\chi_{m}\,\frac{i}{\sqrt{2}}\,\frac{Y}{t^{2}(t^{2}-1)},\qquad q_{3}^{(\ell)}=-\chi_{m}\,i\,\frac{t^{4}+3}{3t^{3}(t^{2}-1)}.$
  • $m$ odd ($\psi_{m},\kappa_{m}$).
    $j=1:\$$q_{1}^{(\ell)}=0,\qquad q_{2}^{(\ell)}=0,\qquad q_{3}^{(\ell)}=0.$
    $j=2:\$$q_{1}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y}{t},\qquad q_{2}^{(\ell)}=i\,\kappa_{m}\,\frac{t^{4}+1}{t^{2}},\qquad q_{3}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\kappa_{m}\,\frac{Y(t^{4}+3)}{3t^{3}}.$
    $j=3:\$$q_{1}^{(\ell)}=i\,\psi_{m}\,\frac{t}{t^{2}+1},\qquad q_{2}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Y}{t^{2}+1},\qquad q_{3}^{(\ell)}=i\,\psi_{m}\,\frac{t^{4}+3}{3t(t^{2}+1)}.$
    $j=4:\$$q_{1}^{(\ell)}=i\,\psi_{m}\,\frac{t^{4}+t^{2}+1}{t(t^{2}+1)},\qquad q_{2}^{(\ell)}=-\frac{i}{\sqrt{2}}\,\psi_{m}\,\frac{Y(t^{4}+t^{2}+1)}{t^{2}(t^{2}+1)},$
    $\hskip 45.00006ptq_{3}^{(\ell)}=i\,\psi_{m}\,\frac{(t^{4}+3)(t^{4}+t^{2}+1)}{3t^{3}(t^{2}+1)}.$
@@
In the case $m$ odd and $j=3$, the coefficient $q_{3}^{(\ell)}$ contains no factor of $Y$.
7 Half-period shift identities at $e_{2}=-1$

### v3

@@ 3 Weierstrass uniformization for the square lattice
Let $\wp(u)$, $\wp^{\prime}(u)$ and $\zeta_{W}(u)$ denote the corresponding Weierstrass elliptic and zeta functions; see, e.g., [14, §23], [15, Ch. 20], [16, Ch. 20].
Injective (birational) uniformization .
On the lemniscatic curve $Y^{2}=2(t^{4}+1)$, set
$\wp(u)=x_{W}(t,Y):=\frac{Y+\sqrt{2}+\sqrt{2}\,t^{2}}{Y+\sqrt{2}-\sqrt{2}\,t^{2}},\qquad\frac{1}{2}\,\wp^{\prime}(u)=y_{W}(t,Y):=-\frac{4t\,(Y+\sqrt{2})}{\bigl(Y+\sqrt{2}-\sqrt{2}\,t^{2}\bigr)^{2}}.$ (21)
Then $(x_{W},y_{W})$ satisfy $y_{W}^{2}=x_{W}^{3}-x_{W}$, hence (21) defines a birational isomorphism between $\Sigma_{\mathrm{lem}}$ and the Weierstrass cubic with invariants (20). In particular, the map (21) is injective on $\Sigma_{\mathrm{lem}}$ (it does not identify $(t,Y)$ with (t,Y)(-t,-Y)), which is essential for the scattered-allocation argument in Theorem 11.1.
The physical lift $u=u(\zeta)$ is selected by composing the physical branch $\zeta\mapsto(t(\zeta),Y(\zeta))$ from (19) with the uniformization (21). Let $u_{0}$ denote the half-period corresponding to the physical point $(t,Y)=(0,-\sqrt{2})$, so that
@@ 4 Pole set, labels, and incident exclusion
$(+,+)\mapsto 3,\quad(+,-)\mapsto 1,\quad(-,+)\mapsto 4,\quad(-,-)\mapsto 2,$ (24)
@@
For later reference we make the pole label $\ell\mapsto(t_{\ell},Y_{\ell})\mapsto u_{\ell}$ explicit. Set the limiting-absorption incident angle $\zeta_{i}=\theta_{i}+\mathrm{i}\varepsilon$ and fix $\theta_{w}=\pi/4$. For $(\sigma,\varepsilon_{w})\in\{\pm 1\}^{2}$ define the four forcing phases
$a_{\sigma,\varepsilon_{w}}:=\exp\!\big(\mathrm{i}\sigma(\zeta_{i}+\varepsilon_{w}\theta_{w})\big),\qquad b_{m,\sigma,\varepsilon_{w}}:=a_{\sigma,\varepsilon_{w}}^{(-1)^{m}}\quad(m=0,1,2,3).$ (27)
The physical orbit table fixes the pole condition in the form $e^{\mathrm{i}w_{m}}=(s(\tau^{m}p))^{(-1)^{m}}=b_{m,\sigma,\varepsilon_{w}}$, so that the forcing pole points are solutions of $s(q)=b$ on $\Sigma_{\mathrm{lem}}$ with $q=\tau^{m}p$.
@@
On the physical component $\Omega_{\mathrm{phys}}^{+}$ one selects the root $t_{\mathrm{in}}(b)\in\mathbb{D}:=\{|t|<1\}$, and the other root is $t_{\mathrm{out}}(b)=1/t_{\mathrm{in}}(b)$.
@@
$\wp(u_{\ell})=x_{W}(t_{\ell},Y_{\ell}),\qquad\frac{1}{2}\,\wp^{\prime}(u_{\ell})=y_{W}(t_{\ell},Y_{\ell}),$ (32)
with the physical lift selected on $\Omega_{\mathrm{phys}}^{+}$.
@@ 5.1 Derivative and residue conventions
$w_{m}^{\prime}(u):=\frac{\,\mathrm{d}w_{m}}{\,\mathrm{d}u}=\begin{cases}\mathrm{i}\sqrt{2}\Bigl(t-\dfrac{1}{t}\Bigr),&m\ \text{even},\\[6.0pt] -\mathrm{i}\sqrt{2}\Bigl(t+\dfrac{1}{t}\Bigr),&m\ \text{odd},\end{cases}\qquad t=t(u).$ (33)
@@
The orbit derivatives in (33) follow from the lemniscatic Snell relation (16) and the injective uniformization (21).
@@
Write $w=w_{0}$ for the physical branch defined by $e^{\mathrm{i}w}=s(t,Y)$, and set $t=t(u)$, $Y=Y(u)$ along the physical lift. Differentiate $\wp(u)=x_{W}(t,Y)$ using (21) and $Y^{2}=2(t^{4}+1)$:
$\wp^{\prime}(u)=\frac{\,\mathrm{d}}{\,\mathrm{d}u}x_{W}(t(u),Y(u))=\frac{\,\mathrm{d}}{\,\mathrm{d}t}x_{W}(t,Y)\,t^{\prime}(u).$
Since $\wp^{\prime}(u)=2y_{W}(t,Y)$ by (21), a direct algebraic simplification yields
$\frac{\,\mathrm{d}t}{\,\mathrm{d}u}=t^{\prime}(u)=-Y(u).$ (34)
Next, differentiating $\cos w=\nu\cos z$ with $\nu=\sqrt{2}$ and using $t=e^{\mathrm{i}z}$ gives
$\frac{\,\mathrm{d}w}{\,\mathrm{d}z}=g^{\prime}(t,Y)=\frac{\sqrt{2}(t^{2}-1)}{Y},$
which is (17). Since $\mathrm{d}t/\,\mathrm{d}z=\mathrm{i}t$, we have
$\frac{\,\mathrm{d}w}{\,\mathrm{d}t}=\frac{\,\mathrm{d}w/\,\mathrm{d}z}{\,\mathrm{d}t/\,\mathrm{d}z}=-\mathrm{i}\sqrt{2}\,\frac{t^{2}-1}{tY}.$
Combining with (34) gives
$\frac{\,\mathrm{d}w}{\,\mathrm{d}u}=\frac{\,\mathrm{d}w}{\,\mathrm{d}t}\frac{\,\mathrm{d}t}{\,\mathrm{d}u}=-\mathrm{i}\sqrt{2}\frac{t^{2}-1}{tY}\,(-Y)=\mathrm{i}\sqrt{2}\Bigl(t-\frac{1}{t}\Bigr),$
which is the $m$ even case in (33). For the orbit branches $w_{m}$ defined by (15), one has $w_{m}=(-1)^{m}w(\tau^{m}(\cdot))$ (mod $2\pi$), and $g^{\prime}(\tau^{m}(t,Y))=\sqrt{2}(t^{2}-1)/Y$ for $m$ even and $g^{\prime}(\tau^{m}(t,Y))=\sqrt{2}(t^{2}+1)/Y$ for $m$ odd. The additional factor $(-1)^{m}$ for odd $m$ yields the sign in the $m$ odd case of (33). ∎
@@
Equivalently, using (33) and $t=t_{\ell}$,
$r_{I}(\ell)=\begin{cases}-\varepsilon_{j}\,\dfrac{\mathrm{i}\,t_{\ell}}{\sqrt{2}\,\bigl(t_{\ell}^{2}-1\bigr)},&m\ \text{even},\\[10.0pt] \phantom{-}\varepsilon_{j}\,\dfrac{\mathrm{i}\,t_{\ell}}{\sqrt{2}\,\bigl(t_{\ell}^{2}+1\bigr)},&m\ \text{odd}.\end{cases}$ (36)
@@ 5.2 Phase symbols
$\chi_{m}:=\mathrm{i}^{m}\in\{\pm 1\}\ (m\ \text{even}),\qquad\psi_{m}:=\mathrm{i}^{m}\in\{\pm\mathrm{i}\},\qquad\kappa_{m}:=\mathrm{i}^{m+1}\in\{\pm 1\}\ (m\ \text{odd}).$ (37)
@@ 6 Residue data and per-pole jet summands
6.1 Residue data tables for alpha_l, beta_l, and C_l
@@
  • if $m$ is even:
    $\alpha_{\ell}=\begin{cases}-\chi_{m}r_{I}\,t^{2},&j=1,\\ -\chi_{m}r_{I}\,(t^{4}-t^{2}+1),&j=2,\\ 0,&j=3,\\ -\chi_{m}r_{I}\,\dfrac{\mathrm{i}Y}{\sqrt{2}}\,(t^{2}-1),&j=4;\end{cases}$
  • if $m$ is odd:
    $\alpha_{\ell}=\begin{cases}\psi_{m}r_{I}\,t^{4},&j=1,\\ \psi_{m}r_{I},&j=2,\\ \kappa_{m}r_{I}\,\dfrac{Y}{\sqrt{2}}\,t^{2},&j=3,\\ -\kappa_{m}r_{I}\,\dfrac{Y}{\sqrt{2}},&j=4.\end{cases}$
@@
  • if $m$ is even:
    $\beta_{\ell}=\begin{cases}\chi_{m}r_{I}\,\dfrac{\mathrm{i}}{\sqrt{2}}\,Yt^{2},&j=1,\\ \chi_{m}r_{I}\,\dfrac{\mathrm{i}}{\sqrt{2}}\,Y,&j=2,\\ -\chi_{m}r_{I}\,t^{4},&j=3,\\ -\chi_{m}r_{I},&j=4;\end{cases}$
  • if $m$ is odd:
    $\beta_{\ell}=\begin{cases}0,&j=1,\\ \kappa_{m}r_{I}\,\dfrac{1}{\sqrt{2}}\,Y(t^{2}+1),&j=2,\\ \psi_{m}r_{I}\,t^{2},&j=3,\\ -\psi_{m}r_{I}\,(t^{4}+t^{2}+1),&j=4.\end{cases}$
@@
  • if $m$ is even:
    $C_{\ell}=\begin{cases}-\chi_{m}\,\dfrac{r_{I}}{2t^{2}},&j=1,\\[6.0pt] -\dfrac{r_{I}}{2}\,\dfrac{2t^{4}-2t^{2}+1}{t^{4}},&j=2,\ m=0,\\[10.0pt] +\dfrac{r_{I}}{2}\,\dfrac{1}{t^{4}},&j=2,\ m=2,\\[8.0pt] 0,&j=3,\\ 0,&j=4;\end{cases}$
  • if $m$ is odd:
    $C_{\ell}=\begin{cases}0,&j=1,\\ 0,&j=2,\\ \dfrac{r_{I}}{2}\Bigl(1+\kappa_{m}\dfrac{Y}{\sqrt{2}t^{2}}\Bigr),&j=3,\\[8.0pt] -\dfrac{r_{I}}{2t^{2}}\Bigl(1+\kappa_{m}\dfrac{Y}{\sqrt{2}t^{2}}\Bigr),&j=4.\end{cases}$
@@ 6.2 Per-pole jet summands
with coefficients $p_{n}=\sum_{\ell\in I_{\mathrm{scat}}}p_{n}^{(\ell)}$ and $q_{n}=\sum_{\ell\in I_{\mathrm{scat}}}q_{n}^{(\ell)}$. For later reference we record the per-pole contributions to the jet-killing coefficients in closed form. The derivation is given in Appendix A; the only changes are the half-period shift data (Section 7) and the local scale $\delta\sim t/\sqrt{2}$ (Section 8.2).
@@
Evaluate at $(t,Y)=(t_{\ell},Y_{\ell})$ and denote $D:=Y+\sqrt{2}$. Then
$p_{1}^{(\ell)}=\frac{1}{\sqrt{2}}\,\alpha_{\ell}\,W_{0\ell},\qquad p_{2}^{(\ell)}=\frac{1}{4}\,\alpha_{\ell}\,W_{1\ell},\qquad p_{3}^{(\ell)}=\frac{1}{12\sqrt{2}}\,\alpha_{\ell}\,W_{2\ell},$
where $(W_{0\ell},W_{1\ell},W_{2\ell})$ are given in (39).
@@
Evaluate at $(t,Y)=(t_{\ell},Y_{\ell})$ and denote $D:=Y+\sqrt{2}$. Then
$q_{1}^{(\ell)}=\frac{1}{\sqrt{2}}\,\beta_{\ell}\,W_{0\ell},\qquad q_{2}^{(\ell)}=\frac{1}{4}\,\beta_{\ell}\,W_{1\ell},\qquad q_{3}^{(\ell)}=\frac{1}{12\sqrt{2}}\,\beta_{\ell}\,W_{2\ell},$
where $(W_{0\ell},W_{1\ell},W_{2\ell})$ are given in (39).
@@
For fully explicit “one-line” formulas in terms of $t_{\ell},Y_{\ell}$ only, one may substitute the tables in Propositions 6.16.2 and the explicit $r_{I}$ form (36) into Propositions 6.46.5 and simplify using $Y^{2}=2(t^{4}+1)$. We keep the compact factorized form above because it is both verifiable and robust under algebraic refactoring.
7 Half-period shift identities at e2 = -1